Abstract. We show how to find arbitrarily fast convergent series expansions for π of the form π = ∑_{n=0}^{∞} S(n)/\binom{mn}{pn}a^{n}, where S(n) is some polynomial in n (depending on m, p, a). We prove that there exist such expansions for m = 8k, p = 4k, a = (−4)^{k}, for any k, and give explicit examples for such expansions for small values of m, p and a.